# Stable Differentiable Modal Synthesis for Learning Nonlinear Dynamics

### v2

@@ 6.1  Datasets
Simulation parameters used for the datasets are provided in Table 1, where $T_{\mathrm{sim}}$ corresponds to the duration of simulation. For specified parameter ranges, randomised values were generated from a uniform distribution. For training, fundamental frequencies, equal to $\frac{\gamma}{2}$, spanned a half-octave range from $\mathrm{61{.}74}$$\mathrm{Hz}$ (B1 note) to $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note). For validation and testing, the next half-octave range from $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note) to $\mathrm{123{.}47}$$\mathrm{Hz}$ (B2 note) was used. In addition, the simulation duration was increased to account for a longer decay time due to a smaller loss parameter $\sigma_{0}$. Non-overlapping ranges for the stiffness parameter $\kappa$ were used. Strings were excited by randomly-generated excitation functions (2.3) at randomised excitation positions $x_{\mathrm{e}}$. The range for the excitation amplitude $f_{\mathrm{amp}}$ was scaled linearly with fundamental frequencies to preserve the strength of the nonlinear effects between the datasets, and thus the range of displacements $\mathbf{q}$ (see Appendix A.1). An audio output was drawn from randomised positions $x_{\mathrm{o}}$ along a string for each trajectory.
@@ A.1  Adjustment for Excitation Amplitude
We consider the following initial value problem for a harmonic oscillator under the driving function $f_{\mathrm{e}}(t)$ (2.3):

### v3

@@ 6.1  Datasets
Simulation parameters used for the datasets are provided in Table 1, where $T_{\mathrm{sim}}$ corresponds to the duration of simulation. For specified parameter ranges, randomised values were generated from a uniform distribution. For training, fundamental frequencies, equal to $\frac{\gamma}{2}$, spanned a half-octave range from $\mathrm{61{.}74}$$\mathrm{Hz}$ (B1 note) to $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note). For validation and testing, the next half-octave range from $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note) to $\mathrm{123{.}47}$$\mathrm{Hz}$ (B2 note) was used. In addition, the simulation duration was increased to account for a longer decay time due to a smaller loss parameter $\sigma_{0}$. Non-overlapping ranges for the stiffness parameter $\kappa$ were used. Strings were excited by randomly-generated excitation functions (3) at randomised excitation positions $x_{\mathrm{e}}$. The range for the excitation amplitude $f_{\mathrm{amp}}$ was scaled linearly with fundamental frequencies to preserve the strength of the nonlinear effects between the datasets, and thus the range of displacements $\mathbf{q}$ (see Appendix A.1). An audio output was drawn from randomised positions $x_{\mathrm{o}}$ along a string for each trajectory.
@@ A.1  Adjustment for Excitation Amplitude
We consider the following initial value problem for a harmonic oscillator under the driving function $f_{\mathrm{e}}(t)$ (3):